### v2

# Contextualized Messages Boost Graph Representations

Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This success has prompted several studies to explore the representational capability of GNNs based on the graph isomorphism task. These works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few theoretical works study GNNs with uncountable node feature representation. This paper presents a novel perspective on the representational capability of GNNs across all levels – node-level, neighborhood-level, and graph-level – when the space of node feature representation is uncountable. Specifically, it relaxes the injective requirement in previous works by employing an implicit pseudometric distance on the space of input to create a soft-injective function. This allows distinct inputs to produce similar outputs only if the pseudometric deems the inputs to be sufficiently similar on some representation, which is often useful in practice. As a consequence, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes non-linear and contextualized transformation of neighborhood feature representations is proposed. A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution in context, establishing SIR-GCN as a generalization of classical GNN methodologies. Experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others [12, 19]. GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation [14, 47, 48]. This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph [27, 32, 47, 50].
Among the most widely used GNNs in literature include the graph convolution network (GCN) [25], graph sample and aggregate (GraphSAGE) [15], graph attention network (GAT) [7, 40], and graph isomorphism network (GIN) [47] which largely fall under the message-passing neural networks (MPNNs) [14] framework. These models have gained popularity due to their simplicity and remarkable performance across various applications [12, 17, 19, 22, 24, 28]. Improvements are also constantly being proposed to achieve state-of-the-art performance [4, 6, 21, 31, 38, 44, 49].
Notably, these advances are mainly driven by heuristics and empirical results. Nonetheless, several studies have also begun exploring the representational capability of GNNs [2, 4, 5, 11, 13, 34]. Most of these works analyzed GNNs in relation to the graph isomorphism task. Xu et al. [47] was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test [46]. Subsequent works build upon their results by considering extensions to the original 1-WL test. However, these results only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. [11] proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
This paper presents a simple yet novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable. The key idea is to define an implicit pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This idea is explored across all levels – node-level, neighborhood-level, and graph-level. Based on the results, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the non-linear and contextualized transformation of neighborhood feature representations is proposed. The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the advantages of the proposed model. Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN in closing.
2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{N}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{N}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
$\bm{H_{u}}$ =$=\left\{\!\!\left\{\bm{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}$
$\bm{a_{u}}$ $=\textsc{AGG}\left(\bm{H_{u}}\right)$ (1)
$\bm{h^{*}_{u}}$ $=\textsc{COMB}\left(\bm{h_{u}},\bm{a_{u}}\right),$
where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset [47] of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
Related works
When $\mathcal{H}$ is countable, Xu et al. [47] showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
$F\left(\bm{H}\right)=\sum_{\bm{h}\in\bm{H}}f\left(\bm{h}\right)$ (2)
is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size. This result forms the theoretical basis of GIN.
Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable. In this setting, Corso et al. [11] proved that if $\bigoplus$ comprises multiple aggregators (e.g. mean, standard deviation, max, and min), the hash function
$M\left(\bm{H}\right)=\bigoplus_{\bm{h}\in\bm{H}}m\left(\bm{h}\right)$ (3)
produces a unique output for every $\bm{H}$ of bounded size. This finding laid the foundation for the principal neighborhood aggregation (PNA) [11]. Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
3 Soft-injective functions
Theorem 1 presents an alternative to injective functions when the space of node features $\mathcal{H}$ is uncountable. It considers a soft relaxation to injectivity, henceforth soft-injectivity, which is often useful in practice, especially in classification tasks.
Theorem 1.
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
Theorem 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function operating on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bounds $\delta_{2}$ and $\varepsilon_{2}$ assert the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bounds $\delta_{1}$ and $\varepsilon_{1}$ ensure $g_{u}$ maintains the relationship of elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ is then said to be soft-injective. An illustration is provided in Fig. 1.
$\delta$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(a) Embedded feature space $\mathcal{S}$.
$d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)<\varepsilon<d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(b) Input feature space $\mathcal{H}$.
Figure 1: Soft-injective feature map $g_{u}$ with pseudometric $d_{u}$.
It is worth noting that if $d_{u}$ is squared Euclidean, then Theorem 1 becomes trivial. Nevertheless, the result becomes non-trivial for other choices of $d_{u}$. Corollary 1 extends this result for multisets.
Corollary 1.
Let $\mathcal{H}$ be a non-empty set with a pseudometric $D$ on bounded, equinumerous multisets of $\mathcal{H}$ defined as
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (5)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
where
$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h}).$ (7)
3.1 Soft-isomorphic relational graph convolution network
Corollary 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn. 7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Similarly, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the message function [14] of the aggregation strategy that transforms the individual neighborhood features. Meanwhile, the distance function $D_{u}$ may be interpreted as a kernel distance [23].
Definition 1 (Collision).
Let $G$ be a function. If $G\left(\bm{H}^{(1)}\right)=G\left(\bm{H}^{(2)}\right)$ and $\bm{H}^{(1)}\neq\bm{H}^{(2)}$, a collision is said to have occurred.
To illustrate the utility of $D_{u}$, suppose node $u$ has two neighbors $v_{1}$, $v_{2}$. If $d_{u}$ is the squared Euclidean distance, then a corresponding message function $g_{u}$ is linear. Fig. 2(a) presents the contour plot of the corresponding hash function $G_{u}$, highlighting potential issues arising from hash collisions. Specifically, consider two multisets $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$. If $\bm{h_{v_{1}}}^{(1)}+\bm{h_{v_{2}}}^{(1)}=\bm{h_{v_{1}}}^{(2)}+\bm{h_{v_{2}}}^{(2)}$, then a hash collision occurs and $G_{u}$ produces identical aggregated neighborhood features even if $\bm{H_{u}}^{(1)}$ and $\bm{H_{u}}^{(2)}$ are fundamentally dissimilar on some representation for a given task.
[…]
(a) $g_{u}(\bm{h})=\bm{h}$.
[…]
(b) $g_{u}(\bm{h})=-\bm{h}^{2}$.
[…]
(c) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h})$.
[…]
(d) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h}+\bm{1})$.
Figure 2: Hash functions $G_{u}$ under different message functions $g_{u}$.
In general, hash collisions may occur when aggregating neighborhood features due to $\mathcal{H}$ being uncountable. Within the current framework, one may simply encode knowledge about node features into the pseudometric $D_{u}$. As a result, only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled. This also corresponds to a more complex and non-linear message function $g_{u}$. To illustrate, if node features represent a zero-mean score, $d_{u}$ may be defined as the squared Euclidean distance of the squared score. A corresponding hash function $G_{u}$ in Fig. 2(b) may then be used to detect potentially anomalous neighborhoods since hash collisions are meaningful for this task.
It is worth noting, however, that Corollary 1 holds for every node $u\in\mathcal{N}$ independently. Hence, different nodes may correspond to different $D_{u}$ and $G_{u}$. For simplicity, one may consider only a single pseudometric compactly defined as
$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})$ (8)
with a corresponding soft-injective hash function
$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right)$ (9)
for every node $u\in\mathcal{N}$. This approach preserves the interpretation of $G$ as an aggregation or hash function with an underlying pseudometric distance $D$ that guides and controls hash collisions. The integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function guiding how features of the key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$ are to be embedded and transformed based on the features of the query (center) node $\bm{h_{u}}$. This provides additional context to hash collisions and makes the message function and pseudometric anisotropic [12] or adaptive with respect to the query node. Figs. 2(c) and 2(d) demonstrate this idea where the introduction of a bias term, assuming a function of $\bm{h_{u}}$, shifts the contour plot and produces different aggregated neighborhood features for identical neighborhood features. Nevertheless, one may opt to inject stochasticity into the node features to distinguish between nodes with identical features and neighborhood features and imitate having distinct hash functions $G_{u}$ for every node $u$ with high probability [34].
For a graph representation learning problem, the relational message function $g$ may be modeled as a two-layered multi-layer perceptron (MLP), with an implied pseudometric, following the universal approximation theorem [16] to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (10)
where $\sigma$ is a non-linear activation function, $\bm{W_{Q}},\bm{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\bm{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$. Leveraging linearity, the model has a computational complexity of
$\mathcal{O}\left(\left|\mathcal{N}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
which is comparable to GAT. In practice, $\sigma$ may also be replaced with a deep MLP if modeling $g$ as a shallow two-layer MLP is infeasible. Moreover, since $\bm{h_{u}}$ is already encoded in the aggregation strategy, the combination strategy may simply be an activation function.
In essence, the proposed SIR-GCN is an instance of the MPNN framework where, unlike most MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic [7] transformation of the neighborhood features to obtain contextualized messages.
3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for every graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn. 7. While this result holds for every graph $\mathcal{G}$ independently, one may simply assume a single pseudometric $D$ with a corresponding soft-injective graph readout function $R$ for a set of graphs $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ from task $\mathcal{D}$. Nevertheless, one may opt to integrate additional information about the graph context and structure into the graph readout function $R$ to imitate having distinct $R_{\mathcal{G}}$ for every graph $\mathcal{G}$ and enhance its representational capability further. The virtual super node [14] may be used in this regard.
In practice, $R$ may also be modeled with an MLP, with an implied pseudometric, to obtain the soft-isomorphic graph readout function
$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{N}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
where $\textsc{MLP}_{R}$ represents the corresponding feature map of $R$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
4 Mathematical discussion
The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the contribution. The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the proposed model.
4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation (by extension) may be used in place of the sum aggregation. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
$\bm{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\bm{W}\bm{h_{v}}$ (13)
and
$\bm{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\bm{W}\bm{h_{v}}$ (14)
which recovers GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN. [^1] GraphSAGE with LSTM aggregation is not included in this discussion. The difference lies in the isotropic [12] nature of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
4.2 GAT
Moreover, in Brody et al. [7], the attention mechanism of GATv2 is modeled as an MLP given by
$e_{u,v}=\bm{a}^{\top}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (15)
with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\bm{W}\bm{h_{v}}$. While the model is anisotropic in nature, messages are nevertheless only linearly transformed with node $u$ only determining the degree of contribution through the scalar $e_{u,v}$. Meanwhile, SIR-GCN directly works with the unnormalized attention mechanism in Eqn. 15 and allows the features of the query node to dynamically transform messages. Specifically, if $\sigma=\textsc{LeakyReLU}$ and $\bm{W_{R}}=\bm{A}$, one obtains
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{A}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right)$ (16)
which shows Eqn. 15 becoming a contextualized message in the SIR-GCN model. Nevertheless, GAT and GATv2 may be recovered, up to a normalizing constant, with an appropriate choice of $\sigma$.
4.3 GIN
Likewise, within the proposed SIR-GCN model, one may explicitly add a residual connection in the combination strategy to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (17)
where $\textsc{MLP}_{\text{Res}}$ is a learnable residual network. If $\textsc{MLP}_{\text{Res}}(\bm{h})=(1+\epsilon)\cdot\bm{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, and $\bm{W_{R}W_{K}}=\bm{I}$, then
$\bm{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\bm{h_{u}}+\sum_{v\in\mathcal{N}(u)}\bm{h_{v}}$ (18)
is equivalent to a GIN. Hence, SIR-GCN with residual connection encompasses GIN.
4.4 PNA
Furthermore, SIR-GCN and PNA approach the problem of uncountable node features $\mathcal{H}$ differently, with the former using only a single aggregator which holds theoretically for any number of neighbors. Nevertheless, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using a linear message function $m$ which is equivalent to a linear transformation of $\bm{h_{v}}$ with a different bias term for each $u$
$m\left(\bm{h_{v}},\bm{h_{u}}\right)=\bm{W_{K}}\bm{h_{v}}+\bm{W_{Q}}\bm{h_{u}}=\bm{W_{K}}\bm{h_{v}}+\bm{b_{u}}.$ (19)
When using mean, max, or min aggregators, the query node $u$ only contributes to the aggregated neighborhood features through the bias term $\bm{b_{u}}$. Meanwhile, when using normalized moment aggregators, the query node $u$ no longer contributes to the aggregated neighborhood features. This potentially limits the expressivity of PNA since, in contrast to SIR-GCN and as suggested by Brody et al. [7], messages are only linearly and statically [7] transformed by the features of the query node.
4.5 1-WL test
Additionally, in terms of graph isomorphism representational capability, SIR-GCN is comparable to a modified 1-WL test. Suppose $w^{(l)}_{u}$ is the WL node label of node $u$ at the $l$th WL-test iteration. The modified update equation is given by
$w^{(l)}_{u}\leftarrow\text{hash}\left(\left\{\!\!\!\left\{\left[w^{(l-1)}_{v},w^{(l-1)}_{u}\right]:v\in\mathcal{N}(u)\right\}\!\!\!\right\}\right),$ (20)
where the modification lies in concatenating the label of the center node with every element of the multiset before hashing. This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as noted in the previous section. Thus, SIR-GCN inherits the theoretical capabilities (and limitations) of the 1-WL test.
Overall, SIR-GCN is demonstrated to generalize four prominent GNNs in literature – GCN, GraphSAGE, GAT, and GIN – and is thus at least as expressive as they are. Notably, SIR-GCN offers flexibility in two dimensions of GNNs: aggregation strategy and message transformation. It emphasizes summed aggregation with anisotropic and dynamic (i.e. contextualized) message transformation, making it well-suited for heterophilous tasks [8], but remains adaptable to alternative configurations.
In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator, which theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
SumMeanAnisotropicIsotropicGATGINGraphSAGEGCNSIR-GCN
Figure 3: SIR-GCN expressivity.
5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN. To ensure fairness, models not employing advanced architectural design or manually crafted features using domain knowledge are used as primary comparisons.
5.1 Synthetic datasets
DictionaryLookup
DictionaryLookup [7] consists of bipartite graphs with $2n$ nodes – $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the value of query nodes by matching their attribute with the key nodes. A sample instance is provided in Fig. 4.
Table 1: Test accuracy on DictionaryLookup.
Model $n=10$ $n=20$ $n=30$ $n=40$ $n=50$
GCN 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00
GraphSAGE 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
GATv2 0.99 $\pm$ 0.03 0.88 $\pm$ 0.18 0.74 $\pm$ 0.28 0.56 $\pm$ 0.37 0.60 $\pm$ 0.40
GIN 0.78 $\pm$ 0.07 0.29 $\pm$ 0.03 0.12 $\pm$ 0.03 0.03 $\pm$ 0.00 0.02 $\pm$ 0.01
SIR-GCN 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
Figure 4: DictionaryLookup.
$A$$B$$A$$B$
Figure 5: GraphHeterophily.
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $n$. SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic nature. However, it is observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic nature. The results underscore the utility of an attentional/relational mechanism in capturing the relationship between the query and key nodes.
GraphHeterophily
GraphHeterophily is an original synthetic dataset. It consists of random directed graphs with each node assigned one of $c$ classes. The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph. A sample instance is provided in Fig. 5.
Table 2: Test mean squared error on GraphHeterophily.
Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
GCN 22749 $\pm$ 1242 50807 $\pm$ 2828 62633 $\pm$ 3491 68965 $\pm$ 3784 72986 $\pm$ 4025
GraphSAGE 22962 $\pm$ 1215 36854 $\pm$ 2330 30552 $\pm$ 1574 21886 $\pm$ 1896 16529 $\pm$ 1589
GATv2 22329 $\pm$ 1307 44972 $\pm$ 2834 49940 $\pm$ 2942 50063 $\pm$ 3407 49661 $\pm$ 3488
GIN 39.620 $\pm$ 2.060 37.193 $\pm$ 1.382 34.649 $\pm$ 1.502 32.424 $\pm$ 1.841 30.091 $\pm$ 1.429
SIR-GCN 0.001 $\pm$ 0.000 0.004 $\pm$ 0.005 1.495 $\pm$ 4.428 0.038 $\pm$ 0.068 0.089 $\pm$ 0.134
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $c$. SIR-GCN achieves an MSE loss of nearly 0 attributed to its anisotropic nature and sum aggregation. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. [47]. Meanwhile, GIN successfully retains the graph structure but fails to learn the relationship between the labels of the query (center) node and key (neighboring) nodes. The results illustrate the utility of anisotropic models even in graph property prediction tasks with countable node features.
5.2 Benchmark datasets
Benchmarking GNNs
Benchmarking GNNs [12] is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. Dwivedi et al. [12] provides more information regarding the individual datasets.
Table 3: Test performance on Benchmarking GNNs.
Model WikiCS ($\uparrow$) PATTERN ($\uparrow$) CLUSTER ($\uparrow$) MNIST ($\uparrow$) CIFAR10 ($\uparrow$) ZINC ($\downarrow$)
MLP 59.45 $\pm$ 2.33 50.52 $\pm$ 0.00 20.97 $\pm$ 0.00 95.34 $\pm$ 0.14 56.34 $\pm$ 0.18 0.706 $\pm$ 0.006
GCN 77.47 $\pm$ 0.85 85.50 $\pm$ 0.05 47.83 $\pm$ 1.51 90.12 $\pm$ 0.15 54.14 $\pm$ 0.39 0.416 $\pm$ 0.006
GraphSAGE 74.77 $\pm$ 0.95 50.52 $\pm$ 0.00 50.45 $\pm$ 0.15 97.31 $\pm$ 0.10 65.77 $\pm$ 0.31 0.468 $\pm$ 0.003
GAT 76.91 $\pm$ 0.82 75.82 $\pm$ 1.82 57.73 $\pm$ 0.32 95.54 $\pm$ 0.21 64.22 $\pm$ 0.46 0.475 $\pm$ 0.007
GIN 75.86 $\pm$ 0.58 85.59 $\pm$ 0.01 58.38 $\pm$ 0.24 96.49 $\pm$ 0.25 55.26 $\pm$ 1.53 0.387 $\pm$ 0.015
GatedGCN - 84.48 $\pm$ 0.12 60.40 $\pm$ 0.42 97.34 $\pm$ 0.14 67.31 $\pm$ 0.31 0.435 $\pm$ 0.011
PNA - - - 97.19 $\pm$ 0.08 70.21 $\pm$ 0.15 0.320 $\pm$ 0.032
EGC-M - - - - 71.03 $\pm$ 0.42 0.281 $\pm$ 0.007
SIR-GCN 78.06 $\pm$ 0.66 85.75 $\pm$ 0.03 63.35 $\pm$ 0.19 97.90 $\pm$ 0.08 71.98 $\pm$ 0.40 0.278 $\pm$ 0.024
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al. [12] to ensure fairness. The results show that SIR-GCN consistently outperforms popular GNNs in literature. Notably, SIR-GCN also outperforms both PNA [11] and efficient graph convolution (EGC-M) [39] which use multiple aggregators. This highlights the significance of non-linear and contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
ogbn-arxiv
ogbn-arxiv [19] is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph [42]. Each node represents an arXiv paper and a directed edge represents a citation. The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
Table 4: Test accuracy on ogbn-arxiv.
Model GIANT-XRT [10] BoT [45] C&S [20] Others Accuracy Parameters
GATv2 0.7415 $\pm$ 0.0005 207,520
GraphSAGE 0.7435 $\pm$ 0.0014 546,344
SIR-GCN 0.7525 $\pm$ 0.0009 667,176
0.7574 $\pm$ 0.0020 697,896
LGGNN [30] 0.7570 $\pm$ 0.0018 1,161,640
DRGAT [1]
KD
0.7633 $\pm$ 0.0008 2,685,527
RevGAT [26]
KD, DCN
0.7636 $\pm$ 0.0013 1,304,912
AGDN [38]
self-KD
0.7637 $\pm$ 0.0011 1,309,760
Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers. Unsurprisingly, however, SIR-GCN performs poorly when compared against complex frameworks utilizing more tricks such as the dynamic evolving initial residual GAT (DRGAT) [1], reversible GAT (RevGAT) [26], and adaptive graph diffusion network (AGDN) [38], albeit with only a small performance difference.
ogbg-molhiv
ogbg-molhiv [19] is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds. Node features contain information regarding the atom while edge features contain information regarding the chemical bond. The task is to predict whether or not the molecules inhibit HIV replication.
Table 5: Test ROC-AUC on ogbg-molhiv.
Model GraphNorm [9] VirtualNode [14] Others ROC-AUC Parameters
GIN
FLAG
0.7748 $\pm$ 0.0096 3,336,306
GIN 0.7773 $\pm$ 0.0129 1,518,901
EGC-M 0.7818 $\pm$ 0.0153 317,265
GCN 0.7883 $\pm$ 0.0100 526,201
PNA 0.7905 $\pm$ 0.0132 326,081
SIR-GCN 0.7721 $\pm$ 0.0110 327,901
0.7981 $\pm$ 0.0062 328,201
GSN [6] 0.7799 $\pm$ 0.0100 3,338,701
GSAT [31] 0.8067 $\pm$ 0.0950 249,602
CIN [4] 0.8094 $\pm$ 0.0057 239,745
Table 5 presents the mean and standard deviation of the test ROC-AUC for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that with only a single GNN layer, SIR-GCN outperforms established models in predicting molecules inhibiting HIV replication, highlighting its expressivity. Nevertheless, SIR-GCN fails to compete with more complex models such as the graph stochastic attention (GSAT) [31] and models incorporating domain knowledge such as the cell isomorphism network (CIN) [4].
6 Conclusion
Overall, the paper provides a novel perspective for creating a powerful GNN when the space of node features is uncountable. The central idea is to use implicit pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This concept is demonstrated at all levels from node features to graph-level features. From the results, a novel SIR-GCN is proposed and shown to generalize classical GNN methodologies. The expressivity of SIR-GCN is then empirically demonstrated with synthetic and benchmark datasets, highlighting its relative superiority in outperforming comparable models. Future studies may explore incorporating SIR-GCN into existing frameworks such as using multiple aggregators [11], reversible GNNs [26], and adaptive graph diffusion networks [38] to achieve state-of-the-art performance.
Acknowledgments and Disclosure of Funding
The work is supported by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821 and 22K19833) and by Kyoto University and Toyota Motor Corporation through the joint project Advanced Mathematical Science for Mobility Society.
Appendix A Proofs
Definition 2 (Pseudometric).
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$.
  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
Definition 3 (Conditionally positive definite kernel [36]).
Let $\mathcal{H}$ be a non-empty set. A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$,
$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)\geq 0,$ (21)
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and
$\sum_{i=1}^{N}c_{i}=0.$ (22)
Lemma 1.
The negative of a pseudometric on $\mathcal{H}$ is a conditionally positive definite kernel on $\mathcal{H}$.
Proof.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. For all $N\in\mathbb{N}$ and $\bm{h}^{(0)},\bm{h}^{(1)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$, by the triangle inequality of $d$,
$-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)$ $\geq-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)$ (23)
$=-\sum_{j=1}^{N}c_{j}\sum_{i=1}^{N}c_{i}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}c_{i}\sum_{j=1}^{N}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)=0.$ (24)
Theorem 2 (Hilbert space representation of conditionally positive definite kernels [3, 35, 36]).
Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\bm{h},\bm{h}\right)=0$ for all $\bm{h}\in\mathcal{H}$. There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(1)}\in\mathcal{H}$,
$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (25)
Proof.
See Schölkopf [36]. ∎
See 1
Proof.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. From Lemma 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
$d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}.$ (26)
For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}$ (27)
$\delta_{2}^{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}<\delta_{1}^{2}$ (28)
$\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (29)
Theorem 3.
Suppose $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
$k\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\dfrac{1}{2}\left[\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(0)}\right)-\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(2)}\right)+\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(0)}\right)\right]$ (30)
is positive definite if and only if $\tilde{k}$ is conditionally positive definite.
Proof.
See Schölkopf [36]. ∎
See 1
Proof.
Let $D$ be a pseudometric on bounded, equinumerous multisets of $\mathcal{H}$ defined as
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (31)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. From Lemma 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})+\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})-2\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}}).$ (32)
Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al. [23]. [^2] If $k$ is also integrally strictly positive definite [37], then the hash function $G$ becomes injective. By the linearity of the inner product, it may be shown that
$D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert,$ (33)
where
$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h})$ (34)
and $g$ is the corresponding feature map of the kernel $k$. For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert<\delta_{1}$ (35)
$\delta_{2}^{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert^{2}<\delta_{1}^{2}$ (36)
$\varepsilon_{2}<D\left(\bm{H_{1}},\bm{H_{2}}\right)<\varepsilon_{1}.$ (37)
Appendix B Experimental set-up
All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (dgl, version 2.1.0+cu121, Apache License 2.0) [43] with PyTorch (torch, version 2.2.1, BSD-3) [33] backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The hyperparameters are chosen based on previous results and heuristics without extensive tuning. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
B.1 Synthetic datasets
DictionaryLookup
Adopting Brody et al. [7], the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
GraphHeterophily
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
B.2 Benchmark datasets
Benchmarking GNNs
The dataset is obtained from dgl (version 2.1.0+cu121, Apache License 2.0) with data splits (training, validation, test) following Dwivedi et al. [12]. In line with Dwivedi et al. [12], all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. [41]. Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW [29] optimizer for over a maximum of 500 epochs with a batch size of 128 (whenever applicable) and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
ogbn-arxiv
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections. Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting. The symmetric mean aggregator is used along with existing tricks in literature. Model training is performed with the AdamW [29] optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
ogbg-molhiv
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer, modified to leverage edge features as described in Appendix C, with 300 hidden units, batch/graph normalization, and residual connections. Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting. The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function. Model training is performed with the AdamW [29] optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
Appendix C SIR-GCN extensions
Denote $\bm{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$. Following the intuition presented in Eqns. 8 and 9, SIR-GCN with residual connection may be modified to leverage edge features to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{E}}\bm{h_{u,v}}+\bm{W_{K}}\bm{h_{v}}\right),$ (38)
where $\bm{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (39)
with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features [18].
Furthermore, one may inject inductive bias into the pseudometric $D$ which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g. stock [17] and fMRI [24] data), $g$ may then be aptly modeled using recurrent-type networks.